The potential energy of particle executing simple hormonic motion 0.2 sec after passing the mean position is 1/4 of its total energy. the period of oscillation is
 

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Potential energy=\dfrac{1}{2}kx^2
Total energy=\dfrac{1}{2}kA^2
Therefore after t=0.2s, x=A/2
Usex=Asin \omega t
Time periodT=\dfrac{2\pi}{\omega}